OpenStudy (anonymous):

derivative of sin11xcos11x?

OpenStudy (anonymous):

sin11xcos11x = (1/2) sin (22x) ( sin11xcos11x )' = ( 22/2 ) cos (22x) = ...

OpenStudy (anonymous):

the answer supposedly is -121sin11xcos11x.. that's why I was really confused

OpenStudy (anonymous):

Where do you get that answer?

OpenStudy (anonymous):

Which seems wrong to my understanding Hmm!

OpenStudy (anonymous):

i know! i worked on this problem for an hour trying to understand what was going on.. i got it from the results provided on the back of the book oh well i'll see tomorrow what the answer was